OpenStudy (anonymous):

Write the equation of the line with the given slope, -5/7and y-intercept, -12. A. y = -(5/7)x - 12 B. y = -(5/7)x + 12 C. y = -12x - (5/7) D. y = -12x + (5/7)

OpenStudy (***[isuru]***):

usually the slandered format of an equation of a straight line is y = mx + b where m ( coefficient of x ) = slope b = intercept now u have got the value of m and b already u just have to substitute them in y =mx + b and get the answer... got it :) ?

OpenStudy (anonymous):

exactly.... y = mx + b slope(m) = -5/7 y intercept (b) = -12 so just sub